rishavraj (rishavraj):

Given tht .. a^2 + 6b = -14 b^2 + 8c = -23 c^2 + 4a = 8 find ab + bc + ca

OpenStudy (luckyness):

Please post your question!~ :)

rishavraj (rishavraj):

@ganeshie8

rishavraj (rishavraj):

@mathmath333 bhai try kr ek baar

rvc (rvc):

mujhe assignment likhe de beta chem ka assignment likh rahi hu bahut kaam hai

rishavraj (rishavraj):

kya gunda banegi ....lol...xD try kr ...mai bhi kr raha hun

rvc (rvc):

masti nahi kar rahi

rishavraj (rishavraj):

mai serious hun yaar......try kr.....honestly

rvc (rvc):

kya hai ans?

rishavraj (rishavraj):

nahi pata -_- .... lol NTSE ka hai .....

rvc (rvc):

ans find kar mujhe ek ans mila but im sure its wrong xD

rishavraj (rishavraj):

bata to sahi

rishavraj (rishavraj):

majak kr rahi ...itni jldi ho gaya -_-

rvc (rvc):

India ki team ko bhula :) tag kar

rishavraj (rishavraj):

krte ja :) ... :P

rishavraj (rishavraj):

apna answer to bata de

rvc (rvc):

ab+bc+ca=3b+4c+2a+1

rvc (rvc):

brb in half an hour @ganeshie8

rishavraj (rishavraj):

sach mein yaar bahut ganda majak tha -_- ..huh really ??? -_-

rishavraj (rishavraj):

@rebeccaxhawaii

ganeshie8 (ganeshie8):

what do we know about a, b, c ? are they integers ?

rvc (rvc):

back :P

rishavraj (rishavraj):

It is just mentioned .... Given tht .. a^2 + 6b = -14 b^2 + 8c = -23 c^2 + 4a = 8 find ab + bc + ca @ganeshie8 previously there was an error ... @rvc *clap clap ...... hurrrrrrrrrrrrrxD

rishavraj (rishavraj):

and one more thing I dont know the answer ....... not having the answer key .... rvc solve krna suru ..... majak mt kariyo -_-

rvc (rvc):

xD (a^2+b^2+c^2)

ganeshie8 (ganeshie8):

is this a multiple choice ?

rvc (rvc):

(a+b+c)^2

rvc (rvc):

haan woh bata pehele mcq hai kya?

rishavraj (rishavraj):

a)26 b)16 c)32 d)25 and yeah sorry for late reply ...... this internet is killing me..... rvc tu ky aganeshie k baatein repeat kr rahi hai ..... :P lol jst messing

rishavraj (rishavraj):

@ganeshie8 bhai kuch pata lage.... I got no clue..... bahut kuch try kr liya -_-

rishavraj (rishavraj):

bhai lol hit and trial s ho gaya....lagta hai NTSE wale chahte nahi the k koi full score kare -_- @ganeshie8 @rvc

rishavraj (rishavraj):

26 answer aaya ....... -_-

rvc (rvc):

how?

rishavraj (rishavraj):

@rvc kya kregi jaan k...sab moh maya hai ...lol xD

rishavraj (rishavraj):

@ganeshie8 u there???? :/

rvc (rvc):

-.-

rishavraj (rishavraj):

hit and trial -_-

rishavraj (rishavraj):

@rvc hmmmm ??

ganeshie8 (ganeshie8):

have you deleted the question ?

rishavraj (rishavraj):

nope.... bt still here it is It is just mentioned .... Given tht .. a^2 + 6b = -14 b^2 + 8c = -23 c^2 + 4a = 8 find ab + bc + ca

rishavraj (rishavraj):

@ganeshie8 bt trust they didnt want students to solve this question -_-

rishavraj (rishavraj):

@hartnn join the party :P

OpenStudy (anonymous):

a b + b c + c a = 26

rishavraj (rishavraj):

yup I got tht bt by hit and trial method.... @robtobey lemme know if theres any valid method

rishavraj (rishavraj):

don't tell me everyone viewing this is question is AFK -_-

rishavraj (rishavraj):

@robtobey u gonna tell ???? :/

OpenStudy (anonymous):

Working on it.

rishavraj (rishavraj):

okkk .... cool ..... xD

OpenStudy (pawanyadav):

a=-2 , b=-3 and c=-4

rishavraj (rishavraj):

@Pawanyadav which method did u use??

OpenStudy (pawanyadav):

I think there is some mistake...wait

rishavraj (rishavraj):

trust me I hav been waiting from last 2 hrs -_-

rishavraj (rishavraj):

btw I got the same values bt by Hit and Trial method @Pawanyadav

OpenStudy (pawanyadav):

After too many manipulation I got (a+2)^2+(b+3)^2+(c+4)^2=0 This is only possible when a=-2,b=-3 and c=-4

OpenStudy (radar):

@Pawanyadav You're good.

rishavraj (rishavraj):

if u can plz show the work......as I told before ...I aint having the answer key....bt got the same values

OpenStudy (pawanyadav):

Is answer 26?

rishavraj (rishavraj):

no idea bt I got te same answer.... and tbh nice job... HATS OFF man @Pawanyadav

OpenStudy (pawanyadav):

Add all three equations.. So you will get.. a^2+4a +b2 +6b +c2+8c+29=0 From a2+4a ,we are getting ideas of (a+2)^2 Similarly with (b+3)^2 and (c+4)^2 And we can break 29=4 + 9 + 16 ..and I got the answer ..that I already told.

rishavraj (rishavraj):

no worries got it :))

OpenStudy (anonymous):

Refer to the attachment from Mathematica v9. The second part is a lengthy human calculation where Mathematica does the intermediate computations.

rishavraj (rishavraj):

@robtobey this internet is literally killing me -_- *I am dead ......

OpenStudy (anonymous):

? Where does it hurt ?

rishavraj (rishavraj):

down there..... I mean at the bottom of the heart :P

rishavraj (rishavraj):

bt really wht u used to solve it -_- damn freaky values

OpenStudy (anonymous):

Well I hope that you recover. Have to bail out now.

rishavraj (rishavraj):

lol yeah...... just keep on praying ... :P lol thanks though ..... wasnt knowing it was gonna take like 2 hrs on here :P

rishavraj (rishavraj):

@mathmath333 u must hav a look on this ...... tbh

rishavraj (rishavraj):

@Vijeya3 @563blackghost

rishavraj (rishavraj):

so thanks to all for their participation :)))

OpenStudy (vijeya3):

Kaunse year ke NTSE ka question tha ye?@rishavraj

rishavraj (rishavraj):

pata nahi yaar subah check kr k bata dunga ....

OpenStudy (vijeya3):

-_- okay

rishavraj (rishavraj):

@Vijeya3 dont worry I wont let u miss me ...lol aise qurstions post krte rahunga :)

OpenStudy (vijeya3):

-_- Cool!Cuz I have to give NTSE this year..so u'll be helping me

rishavraj (rishavraj):

ooo for sure girlll.....any time ....... jst lemme know....

OpenStudy (vijeya3):

:)

rishavraj (rishavraj):

:) ;P >_< >.< -_- <3 </3 >> << ho gaya???

OpenStudy (vijeya3):

-_- yeah

rishavraj (rishavraj):

cool....xD

OpenStudy (mathmath333):

\(\large \color{black}{\begin{align} & (a+b+c)^{2}=a^{2}+b^{2}+c^{2}+2(ab+bc+ca)\hspace{.33em}\\~\\ & -29=a^{2}+b^{2}+c^{2}+2(3b+4c+2a)\hspace{.33em}\\~\\ & (ab+bc+ca)\implies (3b+4c+2a)\hspace{.33em}\\~\\ \end{align}}\) just some similarity

rishavraj (rishavraj):

hmmmm I see